OpenStudy (anonymous):

How do I know what equations to use for projectile motion problems? (Or a really helpful explanation on what I'm looking for to solve projectile motion problems because I just don't understand) From these 3 equations, I don't understand what's being done to change them so they fit x and y components: The underscores are v subscript final and initial, the delta is the triangle, I hope this makes sense and thank you in advance v_f = v_i + a(delta)t x_f = x_i + v_i(t) + 1/2at^2 v_f^2 = v_i^2 + 2a(delta)x

OpenStudy (anonymous):

In proj motion, you can break the problem into two part: horizontal (x) motion and vertical (y) motion. a) Horizontal Motion It is a linear motion with no acceleration. Equation: \(v_x=\text{constant} \) and \(x=v_xt\)

OpenStudy (anonymous):

Thank you for replying, but hmm, I still don't feel like I really understand. I"m going to attach a picture, it's a question with the answers already. If it's possible I'd like you to explain the process in what it is I'm supposed to be doing to solve these questions, I'd like to understand how to do this. Thank you again.

OpenStudy (anonymous):

Forgot to attach the file

OpenStudy (anonymous):

Okay, then i'll continue with b) and explain that pic:

OpenStudy (anonymous):

Thank you very much!

OpenStudy (anonymous):

b) Vertical Motion Gravity affect the motion here, giving vertical acceleration of -g. Equation: \[y=y_0+v_{0y}t-\frac12 gt^2 \\ v_y=v_0-gt\] and optionally: \[v_y^2=v_{0y}^2-2gy\]

OpenStudy (anonymous):

Let'smove to part A now: Given initial velocity v0, theta and g. Asked for th (time to reach max height)

OpenStudy (anonymous):

You have any idea what's the condition for a particle to reach it's max height when thrown?

OpenStudy (anonymous):

Do you mean when the vertical velocity is 0, that's when the max height is reached?

OpenStudy (anonymous):

Yep, correct! Now, we know: - vertical vel is zero (v=0) - want to find time (t) - initial velocity ( \(v_{0y}\) ) which equation in vertical motion should we use then?

OpenStudy (anonymous):

I want to say this one vy=v0−gt (sorry I'm not too sure how to do the subscripts)

OpenStudy (anonymous):

but rearranging it to

OpenStudy (anonymous):

t= vy-v0/g

OpenStudy (anonymous):

Good! But you missed the minus sign.\[t=\frac{v_{0y}-v_y}{g}\] Plug in all what we know to this equation (v0, vy)

OpenStudy (anonymous):

Ah, for g? So \[t = 0 - v _{y}/ -9.8\] I'm not too sure about v_{y}

OpenStudy (anonymous):

Okay, vy is 0 at max height. v0y isn't zero, and for g=9.8 (sice we already put the minus of -9.8 from the beginning in the equations). As for v0y (and v0x), its value is: \[v_{0y}=v_0\sin {\theta} \\ v_{0x}=v_0\cos{\theta} \]

OpenStudy (anonymous):

Sorry, so v\[v _{0y}\] is always equal to \[v _{0y}\sin\]theta which is the same as \[v\]

OpenStudy (anonymous):

Vy

OpenStudy (anonymous):

It goes like this: \[ t=\frac{v_{0y}-v_y}{g} \\ t=\frac{v_0\sin{\theta}-0}{g} \\ t=\frac{v_0\sin{\theta}}{g} \] Let me remark important things here: 1) v0y = v0 sin(theta) 2) vy =0 3) actually, v ISN'T equal to 0 (only stops in vertical motion, but still moving in x direction)

OpenStudy (anonymous):

Earlier, I said "- vertical vel is zero (v=0)" Sorry for this, it should be "- vertical vel is zero (vy=0)"

OpenStudy (anonymous):

Okay, I think I'm slowly getting this and vy is only equal to 0 because we want the velocity at max height in order to find the time it takes to get to max height?

OpenStudy (anonymous):

Very good. You understand the point.

OpenStudy (anonymous):

And that's okay, no need to apologize for your mistakes, I'm already grateful for your time as it is, so thank you. So for Part B, I'm not to sure why the equation in the answer is multiplied by 2. Is it because it's taking into account the time it takes to reach max height and then fall from max height?

OpenStudy (anonymous):

Yes, the time needed to reach max height is equal to time needed to fall from max height. To hit the ground, proj need to reach max height and then fall. Right?

OpenStudy (anonymous):

Yes, so I think I understand the reason for the 2 now

OpenStudy (anonymous):

So for part C, I'm not sure at all why sin(theta) is squared and \[v _{0y}\] as well

OpenStudy (anonymous):

Okay, for c we're dealing with these variables: - v0 - theta - g - and y (or H) Which equation of vertical motion you'll choose?

OpenStudy (anonymous):

and also vy=0 (they said "to reach max height")

OpenStudy (anonymous):

I'm going to go with \[v^2_{y} = v^2_{0y} - 2gy\]

OpenStudy (anonymous):

which would turn into

OpenStudy (anonymous):

\[v^2_{0y}/2g\] and we get rid of v^2y because it's 0, right?

OpenStudy (anonymous):

yep, y=v0y^2/2g And you already know what v0y is.

OpenStudy (anonymous):

yep so that would be \[v^2_{0y}\sin \theta^2\] i think

OpenStudy (anonymous):

yes, you do prove part 3 then :D

OpenStudy (anonymous):

v0y^2=v0^2sin^2(theta)

OpenStudy (anonymous):

Thank you! I understand that a lot more. So with Part D, and I guess B as well, why don't we use cos since wouldn't that be for the horizontal components and R is horizontal?

OpenStudy (anonymous):

We do use cos if we want to find the horizontal distance R R=v0x*tH R=v0*cos(theta)*2*tH But to find tH (time needed to reach max height), we just need to see the vertical movement of the object.

OpenStudy (anonymous):

So is it because we're using the vertical components to find tR for part B that we're sticking with sign?

OpenStudy (anonymous):

sticking with sign? which sign do you mean?

OpenStudy (anonymous):

Sorry, I meant sin

OpenStudy (anonymous):

So for part D, I'm a bit lost again on how that equation is gotten to

OpenStudy (anonymous):

yes, the time needed to reach max height only depend on vertical velocity V0y, hence sin. Nothing to do with horizontal velocity, V0x.

OpenStudy (anonymous):

Part D isn't included in your figure above.

OpenStudy (anonymous):

Oh, I'm so sorry, give me one second

OpenStudy (anonymous):

Sorry, i've to off ffor a whhile

OpenStudy (anonymous):

It's okay, thank you so much for all that you've helped me with so far, I am eternally grateful, and I'm so sorry for taking up so much of your time. Thank you again so much!